OpenStudy (anonymous):

Rewrite the following expressions in simplified expontential form. sqrtx^2y^3

OpenStudy (anonymous):

(2y^3)^1/2

OpenStudy (anonymous):

\[\sqrt{x^2y^3}=(x^2y^3)^{\frac{1}{2}}=xy^{\frac{3}{2}}\]

OpenStudy (anonymous):

oops missed the x